Really short question! Please help me to solve, thank you!
5. The following estimates have been made for an investment project: Initial investment: $450,000 Unit price: $40 Product life: 5 years Sales volumes (in units): ye Unit cost of a commodity: $20 Cost ofcapital: 790 year 2 year 3 year 4 year 5 3000 6000 7000 8000 7000 Calculate the sensitivity (i.e. maximum tolerable unfavorable change as a percentage of the original estimated value) in (a) sales price, (b) unit costs (c) sales volume, (d) initial investment.
Answer #1
CALCULATION OF NPV
YEAR SALE VOLUME PROFIT=Volume*($40-$20) PV@7% PRESENT VALUE OF INFLOWS
1 3000 60000 0.934 56040
2 6000 120000 0.873 104760
3 7000 140000 0.816 114240
4 8000 160000 0.763 122080
5 7000 140000 0.713 99820
TOTAL VALUE OF INFLOWS 496940
LESS: INITIAL INVESTMENT 450000
NPV 46940
a)CALCULATION OF SENSITIVITY AS TO UNIT PRICE
LET UNIT PRICE BE=X
YEAR SALE VOLUME PROFIT=Volume*($X-$20) PV@7%
1 3000 3000X-60000 0.934
2 6000 6000X-120000 0.873
3 7000 7000X-140000 0.816
4 8000 8000X-160000 0.763
5 7000 7000X-140000 0.713
TOTAL VALUE OF INFLOWS 24847X-496940
HERE X WILL BE TO MAKE NPV 0 24847X-496940-450000=0
X=(496940+450000)/24847
38.11083833
SENSITIVITY (40-38.11)/40
0.04725
4.73%
b)CALCULATION OF SENSITIVITY AS TO UNIT COST
LET UNIT COST BE=Y
YEAR SALE VOLUME PROFIT=Volume*($40-$Y) PV@7%
1 3000 120000-3000Y 0.934
2 6000 240000-6000Y 0.873
3 7000 280000-7000Y 0.816
4 8000 320000-8000Y 0.763
5 7000 280000-7000Y 0.713
TOTAL VALUE OF INFLOWS 993880-24847Y
HERE Y WILL BE TO MAKE NPV 0 993880-24847Y-450000=0
Y=(993880-450000)/24847
21.88916167
SENSITIVITY OF COST PRICE (21.89-20)/20
0.0945
9.45%
c)SENSITIVITY AS TO SALES VOLUME
AS THE SALE IS IN THE RATIO OF 3:6:7:8:7
LET SALE BE PROFIT=Volume*($X-$20) PV@7%
1 3X 60X 0.934
2 6X 120X 0.873
3 7X 140X 0.816
4 8X 160X 0.763
5 7X 140X 0.713
TOTAL VALUE OF INFLOWS 496.94X
496.94X-450000=0
X=450000/496.94
905.5419165
YEAR NEW SALE VOLUME PRESENT SALE VOLUME SENSITIVITY AS %
1 2716.62 3000 0.09446 9.45
2 5433.24 6000 0.09446 9.45
3 6338.78 7000 0.09446 9.45
4 7244.32 8000 0.09446 9.45
5 6338.78 7000 0.09446 9.45
c) SENSITIVITY AS TO INITIAL INVESTMENT
YEAR SALE VOLUME PROFIT=Volume*($40-$20) PV@7% PRESENT VALUE OF INFLOWS
1 3000 60000 0.934 56040
2 6000 120000 0.873 104760
3 7000 140000 0.816 114240
4 8000 160000 0.763 122080
5 7000 140000 0.713 99820
TOTAL VALUE OF INFLOWS 496940
LESS: INITIAL INVESTMENT 450000
NPV 46940
SENSITIVITY AS TO INITIAL INVESTMENT 0.104311111
10.43%